# Topological zero modes and bounded modes at smooth domain walls: Exact solutions and dualities

### v2

@@ Conclusions
Concluding, we derived the analytical solutions of a modified Jackiw-Rebbi equation describing zero-energy modes at smooth domain walls between topologically trivial and nontrivial phases in one-dimensional topological insulators and superconductors. This allows for the characterizations of topologically-protected zero-modes in terms of a few length scales: the domain wall width, the decay length, and the oscillation wavelength of the wavefunction. This provides a unified framework to describe physically different regimes, from featureless modes localized at sharp domain walls to nonfeatureless modes with "short" or "long" hair localized at smooth domain walls. We also reveal a universal relation between the bulk gap, decay rate, and oscillation momentum of the zero modes, thus providing a quantifiable and experimentally measurable verification of the bulk-boundary correspondence. Additionally, we uncovered an unexpected duality between the topologically protected zero-modes and Shockley surface modes, highlighting a deeper connection between topological and nontopological states of matter. These findings deepen the understanding of the localization properties of edge modes in topological insulators and Majorana zero modes in topological superconductors and shed some new light on the differences and similarities between topological and nontopological zero modes in these systems.
@@ Acknowledgements.
P. M. thanks Junyong Eom and Krishanu Roy Chowdhury for useful discussions. P. M. is supported by the Japan Science and Technology Agency (JST) of the Ministry of Education, Culture, Sports, Science and Technology (MEXT), JST CREST Grant No. JPMJCR19T2, the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Early-Career Scientists KAKENHI Grant No. JP23K13028 and No. JP20K14375.

### v4

@@ Conclusions
In conclusion, we derived the analytical solutions of a modified Jackiw-Rebbi equation describing zero-energy modes at smooth domain walls between topologically trivial and nontrivial phases in one-dimensional topological insulators and superconductors. This allows for the characterizations of topologically-protected zero-modes in terms of a few length scales: the domain wall width, the decay length, and the oscillation wavelength of the wavefunction. This provides a unified framework to describe physically different regimes, from featureless modes localized at sharp domain walls to nonfeatureless modes with "short" or "long" hair localized at smooth domain walls. We also reveal a universal relation between the bulk gap, decay rate, and oscillation momentum of the zero modes, thus providing a quantifiable and experimentally measurable verification of the bulk-boundary correspondence. Additionally, we uncovered an unexpected duality between the topologically protected zero-modes and Shockley surface modes, highlighting a deeper connection between topological and nontopological states of matter. These findings deepen the understanding of the localization properties of edge modes in topological insulators and Majorana zero modes in topological superconductors and shed some new light on the differences and similarities between topological and nontopological zero modes in these systems.
@@ Acknowledgements.
P. M. thanks Junyong Eom and Krishanu Roy Chowdhury for the useful discussions. P. M. is supported by the Japan Science and Technology Agency (JST) of the Ministry of Education, Culture, Sports, Science and Technology (MEXT), JST CREST Grant No. JPMJCR19T2, the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Early-Career Scientists KAKENHI Grant No. JP23K13028 and No. JP20K14375.